# PLAYER*: Enhancing LLM-based Multi-Agent Communication and Interaction in Murder Mystery Games

### v2

@@ Abstract
We propose PLAYER*, a novel framework that addresses the limitations of existing agent-based approaches built on Large Language Models (LLMs) in handling complex questions and understanding interpersonal relationships in dynamic environments. PLAYER* enhances path planning in Murder Mystery Games (MMGs) using an anytime sampling-based planner and a questioning-driven search framework. By equipping agents with a set of sensors, PLAYER* eliminates the need for pre-defined questions and enables agents to navigate complex social interactions. We additionally make a contribution by introducing a quantifiable evaluation method using multiple-choice questions and present WellPlay, a dataset containing 1,482 question-answer pairs. Experimental results demonstrate PLAYER*’s superiority over existing multi-agent methods, enhancing the generalisability and adaptability of agents in MMGs and paving the way for more effective multi-agent interactions.
PLAYER*: Enhancing LLM-based Multi-Agent Communication and Interaction in Murder Mystery Games
@@ 1 Introduction
Recent advancements in LLMs capable of generating human-like responses have boosted the development of LLM-based agents (Soni et al., 2023; Cherakara et al., 2023). Building on this progress, a series of studies focusing on multi-agent communications have showcased the emergence of social interactions, including cooperation (Li et al., 2024a; FAIR et al., 2022), trust (Xu et al., 2023), deception (Wang et al., 2023), and the spread of information (Park et al., 2023). Despite these advances, building agents that intentionally interact with humans in dynamic environments remains challenging due to several limitations. First, current state-of-the-art approaches struggle with complex questions that lack straightforward answers and have difficulty in understanding interpersonal relationships. As a result, in addition to the game rules, agents often require additional human-written guides tailored to specific scenarios, relying on human context understanding (Xu et al., 2023; Wang et al., 2023). This limitation highlights the need for significant manual efforts and reduces the generalisability of these approaches across different domains and situations. Second, applying traditional Multi-Agent Reinforcement Learning (MARL) methods in such scenarios is challenging because of the difficulties in defining state spaces, action spaces, and rewards. The reward associated with the final decision (e.g., identifying the murderer) often represents the only clear certainty (win or lose) (FAIR et al., 2022; Shi et al., 2024). Third, current evaluation methods are difficult to quantify. These methods rely heavily on manual evaluation to gain a deep understanding, but they are considerably influenced by subjectivity (Xu et al., 2023; Wang et al., 2023).
@@
Figure 1: The murder mystery game unfolds as players take on roles, question suspects based on their character’s objectives, and use deduction to identify the killer, while achieving their ideal role-specific ending. The complete game rules can be found in the appendix A.
To address the limitation of lacking generalisability, we propose a novel framework that efficiently optimises path planning in Murder Mystery Games (MMGs), a strategic game involving both cooperation and competition among 4-12 players through natural language negotiation and tactical coordination (as illustrated in Figure 1). Our approach eliminates the need for pre-defined questions or instructions, allowing agents to play various scripts without extra human-written guides, just like human players. This framework enables agents to navigate complex social interactions and make decisions by considering the types of sensors they need. In the case of MMGs, this involves a set of sensors that detect emotions between players, motivations, and opportunities for actions like murder.
To address the difficulty in defining state spaces and rewards, we designed an anytime sampling-based planner instead of using MARL. Unlike informed graph-based search methods, such as A* (Dechter and Pearl, 1985), our planner, inspired by algorithms like RRT* (Karaman and Frazzoli, 2011) and BIT* (Gammell et al., 2020), samples possible states with detective sensors in real-time. The tree-style search path is constructed and pruned dynamically until the target is reached. For MMGs, we combine sensors with a pruner that focuses on extracting highly suspicious murders to construct a purely questioning-driven search framework. PLAYER* is guided by sensors to efficiently generate high-quality questions to uncover more clues, and uses a pruner to more precisely identify the killer.
To address the evaluation problem, we propose more quantifiable and reproducible evaluation criteria by devising a series of multiple-choice questions. This approach encompasses three types of questions: (1) Objective questions about the shared and individual objectives of each character in the game, (2) Reasoning questions probing the reasoning behind the given responses, and (3) Relations questions that are automatically generated from an existing dataset focusing on character relationships (Zhao et al., 2024c). In summary, we have made the following contributions:
(1) We propose PLAYER*, a framework that enhances the generalisability and adaptability of agents by efficiently optimising path planning in MMGs. (2) We designed an anytime sampling-based planner to dynamically construct and prune the search path, enhancing performance and efficiency. (3) We introduce a quantifiable evaluation method using multiple-choice questions focusing on character objectives, reasoning, and facts, thereby reducing the subjectivity in current evaluation practices. We also constructed and annotated a corresponding dataset, WellPlay, which includes 1,482 QA pairs. [^1] Our code and dataset are available at https://github.com/alickzhu/PLAYER.
2 PLAYER*
@@
In this section, we describe the MMG settings and provide an overview of how PLAYER* works.
@@ 2.1 Problem Setting
Figure 2: Problem Setting.
Figure 3: PLAYER* begins with an introduction round, where each agent provides a brief self-introduction. Following this, agents initialise scenario understanding and begin to search for further information by questioning other agents. We use arrows to identify their questioning actions, with blue arrows representing questions asked by the agent itself and yellow arrows representing questions asked by other agents, which are also remembered by the given agent. Based on the responses they receive, agents update their suspicion state, identified by the black arrow (which decides who to ask), and the corresponding action space, identified by the pink area (which decides what questions to ask). The iterative process of inquiry and refinement continues until the final turn, when agents make decisions aligned with their objectives.
In response to the complexities of social interactions in settings such as MMGs, we have developed an innovative interactive framework tailored for such scenarios. As illustrated in Figure 2, this framework entails the creation of a set of agents $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$ for a game with $N_{a}$ playable characters and a set of victims $\mathcal{V}=\{v_{i}\}_{i=1}^{N_{v}}$. Each agent $a_{i}$ is assigned to a character and initialised with its role background script $C_{i}$, suspicious state $s_{i}=[s_{ik}]_{k=1}^{N_{v}}$, and objectives $\mathbf{o}_{i}=\{o_{ij}\}_{j=1}^{N_{i}}$. The script $C_{i}$ is crafted from the unique viewpoint of that character, framing all relationships and events within the story from $a_{i}$’s perspective. The suspicious state $s_{ik}$ is represented as an $(N_{a}-1)$-dimensional vector that encodes agent $a_{i}$’s beliefs about other players’ suspicions regarding the murder of victim $v_{k}$. The vector entries can be either discrete ($\in{\{0,1}\}$) or continuous values ($\in[0,1]$), reflecting the strength of the suspicions. Unlike traditional planning tasks in which the state space can be accessed directly, $s_{i}$ in our scenario is generated from a language model. We define the possible questions as an action space, allowing the derivation of suspicious states through an LLM-based search. For an illustrative example, refer to Appendix A.3.
As shown in Figure 3, each agent is provided with character-specific information, enabling them to interact with others based on their scripts and historical interactions within the game. Subsequently, agents engage in simulated discussions by asking questions and expressing opinions. To ensure fairness, each agent is limited to asking $m$ questions per round over $n$ rounds. After the discussion, agents must collectively decide and vote on the suspected killer.
2.2 Construction of Agents
To enable agents to think and act like human players, we equip them with the following components: (1) Game Rules. The essential information for both human and AI players. (2) Chain-of-Thought. Facilitating a sequential reasoning process to aid in problem-solving. (3) Sensors. In sociology, interpersonal relationships are often conceptualised within a multidimensional space encompassing factors like time, distance, and various other elements (Latané, 1981; Zhao et al., 2021; Trope and Liberman, 2010). Drawing inspiration from this concept, we propose employing a set of sensors associated with each task. For example, in a murder mystery game, these factors could include emotion, motivation, and suspicion. (4) Memory Retrieval. To address the constraints of LLMs’ finite context window and the performance degradation associated with an increasing number of input tokens (Liu et al., 2023a), we implement this module to store and retrieve narrative scripts and dialog logs generated during gameplay. Specifically, we store the embeddings of all generated dialog logs and each agent’s script in a vector database dedicated to each agent. When an agent encounters a new event requiring action, we use the Faiss library (Douze et al., 2024) to retrieve relevant memories. This process enables the construction of prompts tailored to the predetermined maximum script length and dialog history length. It ensures consistent and coherent interactions among agents while assisting in their strategic planning throughout the game. (5) Self-Reflection. Preserving historical experience by reflecting on the current situation to provide a more detailed, accurate, and profound answer.
2.3 PLAYER* Planning Strategy
@@
To navigate an unknown continuous space defined by natural language, which encompasses the dynamics of relationships, perceptions towards the agent, and hidden secrets, PLAYER* approximates the search domain through sampling, and plans the shortest path to the agent’s objective by prioritising searches based on the quality of potential solutions. As illustrated in Figure 4, this framework is fundamentally composed of two key components:
Figure 4: Search and Approximate. PLAYER* generates questions based on sensor-detected values and identifies the killer more precisely with a prunner. This dynamic question generation allows agents to adjust to the specific circumstances of each game session, ensuring a more focused and effective investigation.
Questioning Strategy in Searching based on Sensors
Given the intricate social complexities inherent to MMGs, our proposed framework provides a robust solution through action generation with sensor components. Unlike prior approaches that required predefined prompts and a multitude of pre-set inquiries (Xu et al., 2023), we generate actions directed towards the targets based on the values detected by those sensors. The task-specific sensors can vary across different tasks, and the only requisite input is a list of sensors. For MMGs, we have designated the input sensors $S$ to be $[S_{Emotion},S_{Motivation},S_{Opportunity}]$: Emotion: This sensor assesses an agent’s disposition towards another character, reflecting one’s willingness to assist the given individual or their intent to uncover as many of their issues as possible. Motivation: This sensor evaluates from one’s perspective whether an individual possesses a motive to be the perpetrator. Opportunity Assessment: This sensor objectively determinies whether a character had the opportunity to commit the crime.
The sensor values are obtained based on the related information retrieved from the script and dialog, represented by the function $f_{sensor}:(C^{\prime}_{i},D_{i})\rightarrow S$, where $C^{\prime}_{i}\subseteq C_{i}$ represents the related scripts and $D_{i}$ denotes the associated dialogue history. Then, for agent $a_{i}$ to be questioned, PLAYER* uses LLMs to sample the question $q$ from the action space given the values of sensors $S_{i}$ and related information $(C^{\prime}_{i},D_{i})$, represented by the function $f_{Q}:(C^{\prime}_{i},D_{i},S_{i})\rightarrow q$.
Action Space Refinement with a Pruner
Following the search step, agents received the response from the questioned agent, as well as conversational exchanges among other agents. Each agent first updates its understanding of interpersonal relationships, thereby adjusting sensor values $S$. Subsequently, it also updated $P_{informative}$ based on the revised information and changes in sensor values. Since the number of inquiries an agent can pose is limited, each question incurs a cost, representing a missed opportunity for exploration. The reward for questioning is the information that can be obtained from the anticipated response. As the number of inquiries directed at an agent increases, the assumed exploration within the world space relevant to that agent also rises. Consequently, the expected reward associated with questioning the same agent diminishes proportionally to the number of inquiries already made. Therefore, it becomes essential to evaluate the probability of obtaining valuable information by persisting in questioning a character who has already been asked many questions, even if they appear highly suspicious. We estimate this probability $P_{informative}(D_{i},|Q_{i}|)$ using the previous dialog $D_{i}$ concerning the agent to be questioned, and the number of inquiries already posed $|Q_{i}|$.
In human gameplay, after an initial round of discussion, players often formulate judgments regarding the scenario and direct their focus towards a subset of highly suspicious individuals. This strategic refinement process is critical for conserving cognitive resources and enhancing the efficacy of inferential reasoning. Emulating the cognitive pattern of human beings, we refine the search space $f_{approximate}:(\mathcal{A},S_{i},P_{informative})\rightarrow s_{i}$: guided by sensors, the agent selects the most suspicious targets (See Lower Part of Figure 4).
We did not impose a hard restriction on the refinement process and allow for flexible adjustment of the suspicious state. For instance, the number of suspects could increase, or an agent might shift suspicion to someone previously overlooked. This is because new information might entirely overturn conclusions drawn in prior rounds, and we did not want to limit that possibility. Then we update the action space based on the suspicious state. This approach prevents premature convergence on a suspect and ensures a comprehensive evaluation of all potential leads. By adopting this protocol, agents systematically converge on the most plausible suspects, thereby expediting the resolution of the mystery and heightening the likelihood of identifying the perpetrator within the designated number of discussion rounds.
Input: Agents $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$, Victims $\mathcal{V}=\{v_{i}\}_{i=1}^{N_{v}}$, Suspicious States $s$, Max Round $n$
@@
2 while current_round < $n$ do
       // Action Generation
@@ 5                   $suspect\_list_{ik}$ = Suspect_Generation($s_{ik}$)
7                         question = Action_Generation($a_{i},a_{j},v_{i}$)
@@
      // Update the suspect list
@@
14                   $s_{ik}$ = Update_Suspicious_State($a_{i},a_{j},v_{i}$)
@@
// Evaluation
18 for i = 1 to $N_{a}$ do
19       Result = Evaluate($a_{i}$)
20
Algorithm 1 Anytime Sampling-based Planning in Multi-agent Interaction
Algorithm 1 outlines a detailed procedure of the game process. While initially designed for MMGs, the framework’s adaptable nature allows it to be readily applied to various games by adjusting the game rules and task-specific sensors, thereby enhancing the gaming experience and offering a structured way for agents to handle complex social dynamics efficiently.

### v3

# Questioning the Unknown: Optimising Multi-Agent Collaboration in Narrative-Driven Games

@@ Abstract
We present Questum, a novel framework for Large Language Model (LLM)-based agents in Murder Mystery Games (MMGs). MMGs pose unique challenges, including undefined state spaces, absent intermediate rewards, and the need for strategic interaction in a continuous language domain. Questum addresses these complexities through a sensor-based representation of agent states, a question-targeting mechanism guided by information gain, and a pruning strategy to refine suspect lists and enhance decision-making efficiency. To enable systematic evaluation, we propose WellPlay, a dataset comprising 1,482 inferential questions across 12 games, categorised into objectives, reasoning, and relationships. Experiments demonstrate Questum’s capacity to achieve superior performance in reasoning accuracy and efficiency compared to existing approaches, while also significantly improving the quality of agent-human interactions in MMGs. This study advances the development of reasoning agents for complex social and interactive scenarios.
Questioning the Unknown: Optimising Multi-Agent Collaboration in Narrative-Driven Games
@@ 1 Introduction
Recent advancements in LLMs capable of generating human-like responses have boosted the development of LLM-based agents (Soni et al., 2023; Cherakara et al., 2023). Building on this progress, a series of studies focusing on multi-agent communications have showcased the emergence of social interactions, including cooperation (Li et al., 2024a; FAIR et al., 2022), trust (Xu et al., 2023a), deception (Wang et al., 2023), and the spread of information (Park et al., 2023). Despite these advances, building agents for Murder Mystery Games (MMGs) that intentionally interact with humans in dynamic environments remains challenging.
@@
Figure 1: The murder mystery game involves players taking on roles, questioning suspects, and using deduction to identify the killer, all while pursuing their character-specific objectives.
As shown in Figure 1, MMGs are strategic games that involve 4-12 players assuming character roles with specific scripts and objectives. Gameplay unfolds through language-based negotiation and tactical coordination over $n$ rounds, with players limited to $m$ questions per round. The ultimate goal is to uncover the truth collaboratively while advancing individual objectives. At the end of the game, players vote on the suspected killer and evaluate their objective completion.
MMGs are script-based games in which each script can only be played once. Unlike traditional games such as Go, which have fixed settings and well-defined action spaces for searching, or Werewolves (e.g., voting results) and Avalon (e.g., mission success), which have intermediate rewards, MMGs lack both well-defined searching spaces and intermediate rewards in decision making.
To address these challenges, we first propose using a discrete space, which is called "sensor" to model each character to character states. Since it can be captured or updated by prompting the natural language description, we are able to explore this search space using generative Question-Answering (what to ask) between agents in real time as a regular planning problem (Karaman and Frazzoli, 2011; Gammell et al., 2020). For the issue of intermediate reward in searching, we design an Information Gain (IG) based heuristic objective which aims to balance existing searching results and the expected reward for further questions. After each action, the agent evaluates whether the current information can update the character states and narrow down the range of suspects. Meanwhile, our method focuses on extracting highly suspicious characters, creating a question-target selection framework guided by IG (who to ask). Compared to previous work, this framework optimises the information collection path, significantly improving system efficiency.
Furthermore, current evaluation methods either focus solely on the final win/loss outcome or rely heavily on manual evaluation, which significantly limits the scope of analysis due to the high-cost (Xu et al., 2023a; Wang et al., 2023; Wu et al., 2024). For instance, Wu et al. (2024) manually designed only 56 inferential questions for 4 selected MMGs. To address the evaluation problem, we constructed a dataset, WellPlay, containing 1,482 inferential questions across 12 MMGs. These questions are categorised into three types: objective, reasoning, and relation. Including the win rate metric used in previous works, this provides a more diverse evaluation framework. We also recruited human players to interact with the agent. The results revealed that previous agents often overfit to agent-vs-agent interaction environments, exhibiting verbose and repetitive dialogues that detracted from the player experience. By integrating our goal-oriented pruner, these issues were significantly mitigated, resulting in more human-like agent behaviour and greatly enhancing player satisfaction.
In summary, we have made the following contributions: (1) We propose Questum, a framework using detective sensors to define and optimise the search in continuous language space. (2) We introduce an IG-based reward to address the lack of intermediate rewards, improving agent behaviour and efficiency. (3) We construct WellPlay, a dataset with 1,482 inferential questions across diverse categories, offering a more robust evaluation framework. (4) By integrating a goal-oriented pruner, we enhance agent-human interaction, reducing verbosity and improving player experience.  [^1] Our code and dataset are available at https://github.com/alickzhu/Questum.
2 Questum
@@
In this section, we describe the MMG settings and provide an overview of how Questum works.
@@ 2.1 Problem Setting
In response to the complexities of social interactions in settings such as MMGs, we have developed an innovative interactive framework tailored to such scenarios. This framework entails the creation of a set of agents $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$ and a set of victims $\mathcal{V}=\{v_{k}\}_{k=1}^{N_{v}}$, where $N_{a}$ and $N_{v}$ denote the numbers of playable characters and victims, respectively. Each agent $a_{i}$ is assigned to a character and initialised with the following:
(1) Role $\mathbf{r}_{i}=\{r_{ik}\}_{k=1}^{N_{v}}$: Whether or not they are the murderer of each victim;
(2) Role background script $C_{i}$: Crafted from the unique viewpoint of $a_{i}$;
(3) Objectives $\mathbf{o}_{i}=\{o_{ij}\}_{j=1}^{N_{o_{i}}}$: A set of $N_{o_{i}}$ goals for agent $a_{i}$.
Game Rules are also provided to Agents as essential information. The detailed rules and procedures are provided in Appendix A.
2.2 Questum  Planning Strategy
Figure 2: Search and Approximate. Questum  generates questions based on character states, selecting agents to question based on past observations of critical information and the likelihood of uncovering more. The goal is to minimise the suspect list.
@@ 2.2 Questum  Planning Strategy
Questum approximates the search domain through sampling, and plans the shortest path to the agent’s objective by prioritising searches based on the quality of potential solutions. As illustrated in Figure 2, this framework is fundamentally composed of two key components: (1) Search based on sensors: Questum  searches for the murderer based on their proximity to the ideal murderer in the language space. (2) Approximation with a pruner: Questum  focuses on a subset of characters that are highly suspicious of being a murderer, and decides whom to query based on this suspect list.
In MMGs, the suspect status is $\mathbf{s_{i}}=\left[s_{ijk}\right]_{j=1,k=1,j\neq i}^{N_{a},N_{v}}$, where $s_{ijk}\in[0,1]$. Here, $s_{ijk}$ encodes agent $a_{i}$’s beliefs about player $a_{j}$’s likelihood of being victim $v_{k}$’s murderer. In our setting, $s_{ijk}$ is set to be discrete ($\in\{0,1\}$), representing suspect and non-suspect. While our list of suspects is finite, our actual search space $L\in\mathbb{R}^{d}$ (the language space) is continuous and high-dimensional, where $d$ is potentially very large. It encompasses the dynamics of relationships, perceptions towards the agent, and hidden secrets.
Search based on sensors
We map each suspect to a point in the state space $L$. Our search process operates continuously in the state space $L$, where character attributes such as emotions, motivations, and opportunities are represented as vectors $u_{i}=(e_{i},m_{i},p_{i},...)$ in this multi-dimensional space. Character attributes are chosen with domain knowledge, drawing on insights from sociology (Latané, 1981; Zhao et al., 2021; Trope and Liberman, 2010). We use sensors to sample the values of these attributes for two main purposes:
1) Estimating the distance between a given suspect and the hypothetical murderer. The hypothetical murderer is assumed to be a point in the search space. At the beginning of the game, when participants lack additional information, the hypothetical murderer refers to an idealised profile that meets specific criteria across all sensor dimensions: someone with motivation, negative emotions, and the opportunity to commit the crime. As the game progresses and more available information, agents update the criteria defining the hypothetical murderer. For instance, if it is suddenly revealed that the victim’s death was accidental, the murderer may no longer be expected to have a motive.
2) Exploration by generating questions. It is modelled as $q=g(u_{i},X)$, where $q$ is a newly generated question by language model $g$. In the inputs, $X$ is the relevant information retrieved from the context by Retrieval-Augmented Generation (RAG) and $u_{i}$ is the state representation of suspect $i$.
Approximation with a pruner
Following the questioning step, agents receive responses from the questioned agent as well as conversational exchanges among other agents. Each agent then updates its understanding of interpersonal relationships, thereby updating character states. In human gameplay, players often form judgments regarding the scenario and focus on a subset of highly suspicious individuals. This strategic refinement process is crucial for conserving cognitive resources and enhancing the efficacy of inferential reasoning. Inspired by the human thinking process, we also enable agents to consider the suspect list.
In an MMG with $N_{a}$ players, to find out who is the murderer $Y$ of a given victim $v_{k}$, we define the entropy to be $H=-ln(\frac{1}{n})$, where $n$ is the number of suspects in the agent’s suspect list. Based on the size of the suspect list for each round of the game, the agent can calculate the corresponding entropy. The target is to maximise this entropy. The initial entropy is $H_{0}=-ln(\frac{1}{N_{a}-1})$, where everyone except the agent itself, who aims to find the murderers, are all suspects, representing the uncertainty about who the murderer is at the start of the game. For each character, we have its historical IG being $IG_{i}=H_{i-1}-H_{i}$ for round $i$, which represents the reduction in uncertainty about the murderer’s identity after receiving the sensor reading and the corresponding answer. $IG_{1:i}=[IG_{1},IG_{2},...,IG_{i}]$. The weighted historical IG is:
$IG_{weighted}[i]=\frac{\sum_{k}w_{k}\times IG_{k}}{\sum_{k}w_{k}}$ (1)
where $w_{k}$ are weights, set to $e^{-(i-k)}$.
After the search step, the agent acquires more information. However, accessing the conditional probability to update the conditional entropy $H_{i+1}=H(Y|u_{i},X)$ might be challenging in closed-source platforms like ChatGPT, and open-source projects such as Llama could introduce strong bias as they usually focus on one suspect Xu et al. (2023b). Hence, we propose a heuristic estimator based on prompting. First, we ask the LLMs to estimate the expected IG. We then convert the discrete value from the prompt to a numerical value:
$E[IG_{i+1}]=\begin{cases}1,&\text{if prompt returns ``high''}\\ 0,&\text{if prompt returns ``medium''}\\ -1,&\text{if prompt returns ``low''}\end{cases}$ (2)
Thus, even closed-source platforms are able to estimate conditional entropy, and the bias in open-source models is reduced by smoothing in the form of multiple-choice questions. We then combine historical and expected gain to determine the highest-scoring character to ask:
$Score_{i}=\beta IG_{weighted}[i]+(1-\beta)E[IG_{i+1}]$ (3)
In addition, we also implement an $\epsilon$-greedy strategy to choose a random character with probability $\epsilon$, and the highest-scoring character with probability $1-\epsilon$ to balance exploration and exploitation. This approach prevents premature convergence on a suspect and ensures a comprehensive evaluation of all potential leads. Thus, agents systematically converge on the most plausible suspects, thereby expediting the resolution of the mystery and increasing the likelihood of identifying the perpetrator within the designated number of discussion rounds. The detailed procedure is in Algorithm 1.
Input: Agents $\mathcal{A}=\{a_{i}\}_{i=1}^{N_{a}}$, Victims $\mathcal{V}=\{v_{k}\}_{k=1}^{N_{v}}$, Suspicious States $s$, Max round $n$.
@@
2 while current_round $\leq$ $n$ do
       // Search based on sensors
@@ 5                   $suspect\_list_{ik}$ = Suspect_Generation($s_{ik}$)
7                         question = Action_Generation($a_{i},a_{j},v_{k}$)
@@
      // Approximation with a pruner
@@
14                   $s_{ik}$ = Update_Suspicious_State($a_{i},a_{j},v_{k}$)
@@
Algorithm 1 Questum Framework